Say Anna's utility function was given by UA = (MA)(MM), where MA is Anna's wealth and MM is Marie's wealth. Initially, Anna has 160 units of wealth and Marie has 40.
In order to maximize her utility Anna should

transfer all her wealth to Marie.

transfer 60 units to Marie.

keep all her wealth for herself.

get 40 units from Marie.
Answer #1

Ans) the correct option is transfer 60 units to Marie.

Anna's current utility is 160*40=6,400

Anna should split their wealth equally: 160+40=200.
200/2=100.
So should give 60 units to Marie. Now Her new utility is 100*100=10,000